OpenStudy (ramennoodle01):

Some friends tell you that they paid $25,000 down on a new house and are to pay $525 per month for 30 yrs. If interest is 7.8% compounded monthly, what was the selling price of the house? How much interest will they pay in 30 years?

OpenStudy (mathmate):

Just like the previous problem, using the same symbols, A=monthly payment P=present value n=number of periods (months) i=nominal annual interest/12 R=1+i Using \(PR^n = A(R^n-1)/i\) you know 4 of the 5 variables, so you can solve for the remainder, whatever is given to you. In this case, to find the selling price, solve for P, and add $25000. Interest is total amount paid, less original amount (P, or present value). = A*n-P

OpenStudy (ramennoodle01):

So A=525 n=(12*30) i= (.078/12) R= 1+ (.078/12)